15. Februar 2002. U. Schoenwaelder; http://www.math.rwth-aachen.de/∼Ulrich.Schoenwaelder HB = Hochschulbibl. RWTH, HBZ = http://www.hbz-nrw.de/ (HBZ-CD-ROM Online), MB = Mathematikbibl., DB = Didaktikbibl. (Winter), EB = Erziehungswiss. Bibl., FH = Bibl. Fachhochschule Aachen, FL = Fernleihe, IB = Informatikbibliothek, IB Nr. = Institutsbibliothek Nr., LB = HB– Lehrbuchsammlung, LS = HB–Lesesaal

Literatur: Kettenbrüche

[1] A. G. Akritas. Elements of Computer Algebra with Applications. J. Wiley & Sons, 1989. MB: 14888. S. 44-52: Euclid’s Algorithm and Continued Fractions.

[2] P. S. Alexandroff, A. I. Markuschewitsch, and A. J. Chintschin. Enzyklopädie der Elementarmathematik, Band I: Arithmetik. Hochschulbücher für Mathematik. Deutscher Verlag der Wissenschaften, 1954. HB: 1 Bb 1054. S. 225–313: Die Elemente der Zahlentheorie (A. J. Chintschin): Kap. III: Der Euklidische Algorithmus und die Kettenbrüche; Kap. IV: Die Darstellung der Zahlen durch systematische Brüche und durch Kettenbrüche; Kap. V: Kettenbrüche und diophantische Approximationen.

[3] G. E. Andrews, B. C. Berndt, L. Jacobsen, and R. L. Lamphere. The Continued Fractions Found in the Unorganized Portions of Ramanujan’s Notebooks. Memoirs vol. 99, number 477. MAA, 1992. MB: 16537.

[4] W. S. Anglin and J. Lambek. The Heritage of Thales. UTM. Springer–Verlag, 1995. MB: 17 671. Part I: History and Philosophy of Mathematics. Part II: Foundations of Mathematics; §2: Natural Numbers (Peano’s Approach), §8: Quaternions, §9: Quaternions Applied to Number Theory (sum of four squares), §14: Continued Fractions, §15: the Fundamental Theorem of Arithmetic (irreduzible und prime Elemente), §16: Linear Diophantine Equations, §17: Quadratic Surds, §18: Pythagorean Triangles and Fermat’s Last Theorem.

[5] H. Appelgate and H. Onishi. The slow continued fraction algorithm via 2 × 2 integer matrices. Amer. Math. Monthly, 90(7):443–455, 1983. MB: Z 42.

[6] B. Artmann. Der Zahlbegriff. Moderne Mathematik in elementarer Darstellung 19. Vandenhoeck & Ruprecht, 1983. HB: Bb1126-19+1. Mit historischen Hinweisen. §2.E: Kettenbrüche.

[7] N. M. Beskin. Fascinating Fractions. Little Math. Library. MIR, 1986. Tel. Rev.: Amer. Math. Monthly 95:5(1988), 477. Continued fractions, π.

[8] A. Beutelspacher and B. Petri. Der goldene Schnitt. BI, 1988. MB: 14687. HB: Bb 1700. Rev.: PM 31:7 (1989), 446. Kapitelüberschriften: Die goldene Spirale und die spira mirabilis; Geometrisches Allerlei; Fibonacci–Zahlen; Kettenbrüche, Ordnung und Chaos; Der Goldene Schnitt in der Natur; Kunst, Poesie, Musik, Witz, Übermuth, Thorheit und Wahnsinn.

[9] Michael N. Bleicher. What’s in a Name? In Anatole Beck, Michael N. Bleicher, and Donald W. Crowe, editors, Excursions into Mathematics, chapter 2, pages 389–475. A K Peters, millenium ed. 2000 edition, 2000. FL: SUB Göttingen. Contains an elementary introduction to fractions (Egyptian fractions, Farey fractions, continued fractions, decimal fractions).

[10] E. Bombieri and A. J. van der Poorten. Continued fractions of algebraic numbers. In W. Bosma and A. van der Poorten, editors, Computational Algebra and Number Theory, Mathematics and Its Applications 325, chapter 11, page ? Kluwer, Dordrecht, 1995.

[11] David M. Bressoud. Factorization and Primality Testing. UTM. Springer–Verlag, 1989. Lst.-B. ISBN 0-387-97040-1. Ch. 10: Continued Fractions (!).

[12] C. Brezinski. History of Continued Fractions and Padé Approximants. Springer Series in Computational Mathematics 12. Springer–Verlag, 1990. MB: 15676.

[13] Bronstein. Taschenbuch der Mathematik. Harry Deutsch. Im Internet unter Leitseite der Hochschulbibliothek der RW-TH Aachen; >Volltexte; >Mathematik, Datenverarbeitung; >Bronstein: Taschenbuch der Mathematik; > Inhalt; >Detailliertes Inhaltsverzeichnis; >Arithmetik; >Kettenbrüche (unter Reelle Zahlen”).

[14] J. W. Bruce, P. J. Giblin, and P. J. Rippon. Microcomputers and Mathematics. Cambridge Univ. Press, 1990. HB: Bm 2340. Euklidischer Algorithmus, Kettenbrüche, Fibonacci-Folge, Eulers Phi–Funktion, Legendre–Symbol, Primzahlen, Primzahltests, π, e (ist irrational: S. 278), Differentialgleichungssysteme.

[15] A. J. Chintschin. Die elemente der zahlentheorie. In P. S. Alexandroff et al., editor, Enzyklopädie der Elementarmathematik, Band I: Arithmetik, Hochschulbücher für Mathematik Band 7, page 379. Berlin: VEB Deutscher Verlag der Wissenschaften, Berlin, ^{2}1965. [HB: Bb1054-1+1 ?] LS B0100. Übers. a. d. Russ. S. 273–285: Kap. IV Die Darstellung der Zahlen durch systematische Brüche und durch Kettenbrüche; S. 285–302: Kap. V Kettenbrüche und diophantische Approximation.

[16] H. Cohen. A Course in Computational Algebraic Number Theory. GTM 138. Springer–Verlag, 1993. MB: 16831; HB: Bb 1244-138+1. S. 21-24: Continued Fraction Expansions of Real Numbers.

[17] B. Conrey. Remarks about continued fractions and real quadratic fields. Vortragsankündigung in Notices AMS 43:4 (1996), 512 (Special Session on Number Theory III), Abstract 910-11-107, 1996.

[18] J. H. Conway and R. K. Guy. The Book of Numbers. Springer–Verlag, 1996. MB: 17923. S. 17: How numbers are written; S. 30: Square numbers; S. 33: Triangular numbers; polygonal numbers; tetrahedral numbers; sums of cubes; S. 68: Pascal’s triangle; S. 91: Bell numbers, Stirlilng numbers, Catalan numbers, Bernoulli numbers, Fibonacci numbers (sumflower) [see also S. 202]; S. 127: Primes; S. 146: Sums of two squares; S. 152: Farey fractions and Ford circles; S. 157: Fractions cycle into decimals; S. 171: Pythagorean fractions; S. 176: Continued fractions; S. 181: Geometric problems and algebraic numbers; S. 211: Complex numbers; Gaussian primes; Eisenstein primes; S. 230: Hamilton’s quaternions; S. 237: Some transcendental numbers: π, Liouville’s number, e, ...; S. 265: Infinite and infinitesimal numbers; S. 283: Surreal numbers; games.

[19] H. S. M. Coxeter. Introduction to Geometry. J. Wiley & Sons, 1961. MB: 1898; MB: 15469 (2nd ed. 1969, Wiley Classics Library edition 1989). S. 210: Farey series.

[20] H. Davenport. The Higher Arithmetic. An Introduction to the Theory of Numbers. Hutchinson & Co. Ltd.; Cambridge Univ. Press, 1952, ^{6}1992. MB: 16 723. Inhalt: I Factorization and primes. II Congruences. III Quadratic residues. IV Continued fractions. V Sums of squares. VI Quadratic forms. VII Some Diophantine equations. VIII Computers and the theory of numbers (7. The RSA cryptographic method).

[21] William Derrick and Jack Eidswick. Continued fractions, Chebychev polynomials, and chaos. Amer. Mathematical Monthly, 102(4):337–344, 1995. MB: Z 42.

[22] F. W. Dustmann. ABAKUS Angewandte Mathematik, Materialien für den Unterricht im Differenzierungsbereich. Schöningh, 1995. 1 Navigationsprobleme, 2 Darstellende Geometrie, 3 Optimierung, 4 Parabeln und verwandte Kurven, 5 Iterative Verfahren (5.2 Endliche Kettenbrüche: Planetarium von Huygens, Kalenderrechnung, Astronomie; 5.3 Unendliche Kettenbrüche: Heron–Verfahren), 6 Geometrie auf der Kugeloberfläche, 7 Mathematische Kartenkunde.

[23] A. Engel. Exploring Mathematics with Your Computer. New Mathematical Library 35. MAA, 1993. ISBN 0-88385-636-0. FL: UB Trier 55BA/ENGE/C12317; UB Bielefeld 100/3099875+1 (ohne Diskette). For high school students, undergraduates, teachers. Chapter 7: 62 Continued fractions.

[24] W. N. Everitt. The two–squares theorem of Fermat. Vortrag im Grauiertenkolleg Aachen 20.5.1996, 1996. Every prime number p of the form 4r + 1 is expressible as the sum of two integer squares. Beweis nach Henry J. S. Smith unter Verwendung von Kettenbrüchen (Euklidischer Algorithmus), konstruktiv.

[25] N. I. Fel’dman and Yu. V. Nesterenko. Transcendental Numbers. Encyclopaedia of Mathematical Sciences, A. N. Parshin and I. R. Shafarevich (Eds.), Vol. 44: Number Theory IV. Springer, 1998. MB: 15893d. Ch. 1: Approximation of algebraic numbers; 1.4 Continued fractions.

[26] W. Franz. Kettenbruchdarstellungen von Quadratwurzeln. Praxis der Mathematik, 8:72–75, 1966. HB: Z 1757.

[27] W. Franz. Kettenbrüche für einige transzendente Funktionen und e. Praxis der Mathematik, 9:271–275, 1967. MB: Z 101.

[28] A. Fricke. Quadratische Gleichungen und ihre iterative Lösung. Praxis der Mathematik, 26(1):3–12, 1984. HB: Z 1757. MB: Z 101. Auch Kettenbrüche.

[29] Midhat J. Gazalé. Number: From Ahmes to Cantor. Princeton Univ. Press, 2000. ISBN 0-691-00515-X. HBZ. Inhalt: 1 The Genesis of the Number System. 2 Positional Number Systems. 3 Divisibility and Number System (The Fundamental Theorem of Arithmetic). 4 Real Numbers (.. Pyth. Triples ..). 5 Continued Fractions [S. 185-198]. 6 Cleavages (.. Mediant ..) [S. 199-256]. 7 Infinity.

[30] K. Girstmair. Periodische Dezimalbrüche – was nicht jeder darüber weiß. In A. Beutelspacher et al., editor, Jahrbuch Überblicke Mathematik 1995, pages 163–179. Vieweg, 1995. HB: 1995 Bb 1303. §4: Rekonstruktion von z/n aus wenigen Ziffern.

[31] Hoffmann. P 848. Kettenbruchmethode in der Diophantik. Praxis der Mathematik, 26(12):382–383, 1984. HB: Z 1757. MB: Z 101. Man bestimme die absolut kleinste Lösung der Pell–Gleichung x^{2} − 94y^{2} = 1 mittels der periodischen Kettenbruchentwicklung \sqrt{(}94) =... durch ein möglichst übersichtliches und leicht merkbares Rechenschema.

[32] M. C. Irwin. Geometry of continued fractions. Amer. Math. Monthly, 96(8):696–703, 1989. MB: Z 42.

[33] T. H. Jackson. From Number Theory to Secret Codes. Adam Hilger, 1987. Continued fractions and rational approximants.

[34] L. Jacobson, editor. Analytic Theory of Continued Frractions III. Seminar–Workshop, Redstone 1988, LNM 1406. Springer–Verlag, 1989. MB: 14999.

[35] Thomas Jahnke. Warum heißt der Mittelwertsatz Mittelwertsatz? Der Mathematikunterricht, 1989:207–210, 1989. HB: Bb1256-1989. Die lokale Änderungsrate schränkt die mittlere Änderungsrate ein. Mediant von mittleren Änderungsraten.

[36] W. B. Jones and W. J. Thron. Continued Fractions: Analytic Theory and Applications. Encyclopedia of Mathematics and its Applications 11. Addison–Wesley. Now distr. by Cambridge Univ. Press, 1980. ISBN 0-201-13510-8. HB: Bb1379-11+1.

[37] A. Khintchine. Kettenbrüche. Teubner, Leipzig, 1956. MB: 559.

[38] S. Lang. Introduction to Diophantine Approximations. Springer–Verlag, ^{2}1995. MB: 3771. HB: Bf 5199+2. Appendix B: Continued Fractions for Some Algebraic Numbers. Appendix C: Addendum to Continued Fractions for Some Algebraic Numbers.

[39] L. J. Lange. An elegant continued fraction for π. Amer. Math. Monthly, 106(5):456–458, 1999. MB: Z 42.

[40] A. Leutbecher. Zahlentheorie. Eine Einführung in die Algebra. Grundwissen Mathematik, Springer–Lehrbuch. Springer– Verlag, 1996. MB: 17917. Ch. 6: Gewöhnliche Kettenbrüche.

[41] W. J. LeVeque. Elementary Theory of Numbers. Addison–Wesley Series in Introductory Mathematics. Addison–Wesley, 1962. MB: 1716. Ch. 5: Continued fractions. Ch. 7-6: Pell’s equation and continued fractions.

[42] W. J. LeVeque. Fundamentals of Number Theory. Addison–Wesley, 1977. MB: 9885. Ch. 9.3: Infinite continued fractions; Ch. 9.5: Appplications to Pells’s equation and to factorization; Ch. 9.6: Equivalence of numbers.

[43] L. Lorentzen and H. Waadeland. Continued Fractions with Applications. Studies in Computational Mathematics 3. Elsevier Science (North Holland), 1992. HB: Bf 9399-3. Rekl. Notices AMS 42:11 (1995), inside back cover.

[44] H. Lüneburg. Kleine Fibel der Arithmetik. B. I.–Wissenschaftsverlag, 1987. Rev.: PM 31:1 (1989), 62. Lagrangescher Kettenbruchalgorithmus – Kettenbruchsituation.

[45] Andreas Meisner. Erfahrungen mit einer unterrichtseinheit zum goldenen schnitt – forderungen für ein ausbildungskonzept. In Rolf Biehler, H. W. Heymann, and B. Winkelmann, editors, Mathematik allgemeinbildend unterrichten: Impulse für Lehrerbildung und Schule, IDM–Reihe Untersuchungen zum Mathematikunterricht 21, pages 92–102. Aulis Verlag Deubner, ^{2}1996, 1995. ISBN 3-7614-1798-5. HB: Kb1467-21+2. Goldener Schnitt, Kettenbruchentwicklung, quadratische Gleichungen via Kettenbruch.

[46] M. Neubrand. Kettenbrüche: Beste Näherungen, transzendente Zahlen. Der Mathematikunterricht, ?(5):30–47, 1984. HB: Z 5577. Geschichte: Ch. Huygens.

[47] M. Neubrand. The planetarium of Christiaan Huygens at Leiden and continued fractions. In J. de Lange, editor, Mathématiques pour tous à l’âge de l’ordinateur, Proc. CIEAEM (Commission internationale pour l’étude et amélioration de l’enseignement des matheématiques), Leiden, 1985, pages 379–381. OW & OC Utrecht, 1986. FL: UB Bielefeld 100/3189444+1.

[48] I. Niven and H. S. Zuckerman. An Introduction to the Theory of Numbers. ^{3}1972. MB: 6502.

[49] I. Niven, H. S. Zuckerman, and H. L. Montgomery. An Introduction to the Theory of Numbers. Wiley, 1991. HB: BO 1304 Aufsicht Lesesaal. Farey Fractions and Irrational Numbers; Simple Continued Fractions.

[50] N.N. Continued fractions. http://www.calvin.edu/academic/math/confrac/, Last visited March 1998. Thesis for Mathematics Honor’s Degree, undergraduate level. Chapters: Introduction, History, Theory, Applications, Bibliography and Sources (with web links).

[51] University of Surrey: Department of Computing. An introduction to continued fractions. www.mcs.surrey.ac.uk/Personal/R.Knott/ Fibonacci/cfINTRO.html, visited May 4 1998.

[52] C. D. Olds. Continued fractions. The New Mathematical Library. MAA, 1963.

[53] C. D. Olds. The simple continued fraction expansion of e. Amer. Math. Monthly, 77:968–974, 1970. MB: Z 42. Winner of the Chauvenet Prize for expository writing in mathematics.

[54] O. Perron. Die Lehre von den Kettenbrüchen. Chelsea, 1950. MB: 428.

[55] O. Perron. Die Lehre von den Kettenbrüchen II. Teubner, 1957. MB:.

[56] O. Perron. Irrationalzahlen. Göschens Lehrbücherei 1. de Gruyter, ^{4}1960. MB: 2826. 4. Kap.: Verschiedene Darstellungsformen irrationaler Zahlen (Kettenbrüche); 5. Kap.: Approximation irrationaler Zahlen durch rationale (§37: Gleichzeitige Approximation mehrerer Zahlen).

[57] H. Rademacher. Higher Mathematics from an Elementary Point of View. Birkhäuser, 1982. Ch. 4: Order of Fractions; Ch. 7: On the Approximation of Irrational Numbers by Rational Numbers; Ch. 8: The Ford Circles.

[58] I. Richards. Continued fractions without tears. Mathematics Magazine, 54(4):163–171, 1981. MB: Z 167.

[59] H. Riesel. Prime Numbers and Computer Methods for Factorization. Birkhäuser, 1985, ^{2}1994. MB: 12884. Appendix 8: Continued fractions.

[60] A. M. Rockett and P. Szusz. Continued Fractions. World Scientific, 1992. Rev.: Bull. AMS 30:1 (1994), 109–111.

[61] G. Scheja and U. Storch. Lehrbuch der Algebra - unter Einschluß der linearen Algebra. Teil 3. B. G. Teubner, Stuttgart, 1981. MB: 10 763 c. Kap. II.C Kettenbrüche.

[62] G. Scheja and U. Storch. Lehrbuch der Algebra - unter Einschluß der linearen Algebra, Teil 1. Mathematische Leitfäden. Stuttgart: Teubner, ^{2}1994. ISBN 3-519-12203-0. S. 128–153: Kap. II.C Kettenbrüche.

[63] Rudolf Schmidt and Richard Stender. Aus der Welt der Zahlen. Schriftenreihe zur Gestaltung des mathematischen Unterrichts 7. Otto Salle, 1954. HBZ. Kettenbruchverfahren zur Berechnung von Quadratwurzeln. Nicht für Mittelstufe, aber Extrastunde.

[64] B. Schuppar. Gute rationale Näherungen für reelle Zahlen. Praxis der Mathematik, 31(2):70–89, 1989. HB: Z 1757, MB: Z 101. Anwendung: Periodizität von Sonnen– und Mondfinsternissen (mit einem Literaturhinweis).

[65] M. Shrader-Frechette. Modified Farey sequences and continued fractions. Mathematics Magazine, 54(2):60–63, 1981. MB: Z 167.

[66] W. Sierpiński. Elementary Theory of Numbers. North–Holland, 1988. MB: 14176. Umfassend. Ch. VIII: Continued fractions.

[67] M. A. Stern. Zur Theorie der periodischen Kettenbrüche. J. reine angew. Mathematik, 53(Heft 1):1–102, 1857. MB: Z 79; HB: Z 2025.

[68] U. Storch and H. Wiebe. Lehrbuch der Mathematik – für Mathematiker, Informatiker und Physiker, Band I: Analysis einer Veränderlichen. BI Wissenschaftsverlag, 1989. HB: Bb 1705-1+1. S. 99–102: Kettenbrüche.

[69] H. S. Wall. Analytic Theory of Continued Fractions. Van Nostrand, 1948.

[70] H.-G. Weigand. Kettenbrüche: eine vergessene Insel in der Darstellungswelt reeller Zahlen. mathematik lehren, 87:52–56, April 1998.

[71] Chee Keng Yap. Fundamental Problems of Algorithmic Algebra. Oxford Univ. Press, 2000. ISBN 0-19-512516-9 (cloth). Ch. 14: Continued Fractions. MB: 19032.